### v1

# Reinforcing VLMs to Use Tools for Detailed Visual Reasoning Under Resource Constraints

Abstract
Despite tremendous recent advances in large model reasoning ability, vision-language models (VLMs) still struggle with detailed visual reasoning, especially when compute resources are limited. To address this challenge, we draw inspiration from methods like Deepseek-r1 for VLMs and train smaller-scale models with Group Relative Policy Optimization (GRPO) to use external tools such as zoom. The greatest benefit is obtained with a combination of GRPO learning, a simple reward structure, a simplified tool-calling interface, allocating additional tokens to the result of the tool call, and a training data mix that over-represents visually difficult examples. Compared to similarly-sized baseline models, our method achieves better performance on some visual question-answering (VQA) tasks, thanks to the detailed visual information gathered from the external tool.
1 Introduction
Recent advances in vision-language models (VLMs) have significantly improved performance across a spectrum of multimodal tasks. However, despite the superb capabilities of state-of-the-art models, VLMs suffer from notable limitations in processing fine-grained visual details [18]. Recent studies highlight that VLMs often struggle with tasks requiring detailed visual understanding, where the subtle visual details that humans can easily interpret are neglected by VLMs [24]. This issue is further amplified with high-resolution image inputs [29, 27].
To overcome this challenge, previous research has explored the incorporation of visual search and multiresolution processing techniques into VLM-based systems. Although visual search systems provide additional information on keypoints in images in the VLM, they can suffer from error propagation due to their complex system design [29, 20]. Multiresolution processing methods, on the other hand, add extra visual tokens which represent the crops of the original input image at different resolutions [23]. While those methods may confer a performance boost, they also introduce computational overhead due to additional image tokens. And the multiresolution image tokens themselves still represent general information about the image, and not necessarily the pertinent details of the specific region of interest (ROI) in the image that would be most helpful for the task.
In this paper, we take inspiration from recent demonstrations of tool use in large language models (LLMs) via reinforcement learning (RL) and extend this approach to multimodal VLMs with a particular focus on efficient use of limited resources. In particular, we efficiently train smaller VLMs via GRPO to capture visual details from task-specific ROIs in real-world images by appropriately calling a zoom-in tool. We choose an RL approach also due to a lack of existing data with annotated tool-use trajectory traces that could act as ground truth for supervised fine-tuning of a vision model for detailed visual reasoning.
Our contributions are as follows:
  • We propose a recipe for extending GRPO to tool use in the visual domain with efficiency under constrained resources.
  • We establish a set of parameters that enable a small VLM to learn tool use for visual detail understanding under resource constraints; and we find that the structure of the reward function and the data mix significantly impact success.
  • We present experimental results showing improvement on small-model SOTA on high-resolution VQA datasets.
  • We discuss how the choice of external tool interface, including tokens allocated to the representation of the results of external tool calls, affects performance.
Our ideas and results were developed independently and simultaneously with those reported by [36], which corroborate some of our findings. Although both works describe a method to train a VLM to zoom in, our work uniquely focuses on overcoming the challenges introduced in the small resource regime.
2 Method
Figure 1: In contrast to direct prompting where the model fails to capture the visual details in high-resolution images, RL incentivizes the VLM to use a zoom-in tool to get extra information from the specific ROI to answer vision-oriented questions correctly (bottom). Resizing the image makes the training efficient while preserving the visual details to be noticeable to the VLM.
Our approach is to teach a VLM to use external tools to enhance detailed image understanding via an agentic take on reinforcement learning. In particular, we train a small Qwen2.5-VL-3B-Instruct [^1] All references to the Qwen2.5-VL models in this paper, whether ours or baselines, refer to the instruction-tuned version of these models. model with Group Relative Policy Optimization (GRPO) [6, 19] to use a zoom tool in order to find relevant information in a curated set of training examples. At the time of writing, this model is among the smallest state of the art VLMs. We operate under a limited computational and memory budget of four A100 80 GPUs. Therefore, our method emphasizes efficient use of limited resources for maximum visual understanding. This necessitates limiting training image resolution, the number of tool calls, and an easy-to-learn tool interface. Note that our method does not rely on ground truth trajectories of tool use, since none are generally available for visual detail data. Instead, given the option of calling on a tool and a reward signal, the model learns how to effectively use the tool to solve the problem.
2.1 Efficient tool use under limited resources
Image resolution
In order to limit the memory footprint during training, we downsize images such that the long side does not exceed 1024 pixels. Note that when using the zoom tool the model only has access to this resized image and not the original higher-resolution one. To make sure the model is able to capture the details in the zoomed-in image, we further upscale the zoomed-in crop to the same size as the downsized input image. We find that this strategy helps the model generalize well on high-resolution visual tasks at inference time, which will be discussed in the following sections. In contrast, during evaluation we do not resize images: the zoom-in tool is directly applied to the original high-resolution image to give the model more visual detail.
Limited tool use
Additionally, to reduce memory load, we limit the number of zoom-in tool use calls to 1 at both training and inference times. This limits our approach, as implemented, to tasks that require at most one detailed ROI per image. We enforce this constraint via structured decoding. When sampling responses from the model at both RL training and inference time, we use the regular expression ([^<]*)</think>([^<]*)<answer>([^<]*)</answer> to guide the decoding of the model’s second response. This both prevents the model from invoking the tool, and ensures that there is an <answer> field in the response and an end to the conversation. Limiting the number of tool calls prevents out-of-memory errors from multiple sequences of image tokens appended to the conversation.
Easy-to-learn tool interface
Figure 2: Our YAML-like interface for tool-use calls uses fewer tokens than standard JSON, and is easier for small models to format correctly.
We find that under limited resource constraints, an easy-to-learn tool interface is essential. Our preliminary experiments were performed with a tool invoked with a complex format (JSON) accepting a bounding box in pixel coordinates as a parameter. We found that this approach did not produce any improvements over the baseline model with no tools. The model learned to zoom into very small and often meaningless regions of the image. We hypothesized that a simpler format would be easier to learn. We achieved this in two ways. First, we adopted a YAML-like tool use interface that is less brittle and more concise than JSON (Figure 2). This format contains fewer tokens, helping the model learn tool use more efficiently. Additionally, we changed the tool to accept a keypoint defined in pixel coordinates instead of a bounding box. The tool then returns a fixed crop centered at the keypoint. This allows for the model to be less precise in its localization and increases the likelihood that the tool’s output contains information that is useful to the query.
2.2 Training
Targeting model efficiency, we train a small Qwen2.5-VL-3B-Instruct model with GRPO on a subset of the training data from the TextVQA [22] dataset. This task was chosen as representative of the sort of images that require zooming in on a particular detail in order to find the answer: in this case, to read the characters in a specific location in the image. Note that the model is not exposed to any high-resolution images during the training stage, but our approach helps the model generalize well on such visual tasks at inference time. The system prompt gives the model the option of calling a zoom tool anchored at a point of interest in the image with a fixed bounding box.
Reward shaping
The rewards for a successful visual reasoning trace are sparse as the answer follows from multiple reasoning steps with potential calls to external tools. Therefore, following existing practice, we incentivize what we believe to be the correct system behavior by providing a structured reward: we reward the model for the correct answer, separately for proper formatting, and separately again for successfully invoking the external zoom tool. In particular, $R=\alpha R_{c}+\beta R_{f}+\gamma R_{t}$, where the correctness reward $R_{c}=\lambda R_{a}+(1-\lambda)R_{e}$ is a combination of a reward for the exact correct answer and a soft reward based on the average edit distance from the top three closest human answers; $R_{f}$ is the format reward; and $R_{t}=\text{\# successful tool calls }/\text{ \# attempted tool calls}$. Our most successful run had $\alpha=1$, $\beta=1$, $\gamma=0.1$ and $\lambda=0.5$.
Data mix
We train our model on a curated mix of images from datasets where zooming in to see visual detail is beneficial. We find that the nature of the training data matters in obtaining the best results when it comes to teaching VLMs tool use. It is important to prioritize training on those images and tasks that are most difficult for the base model without access to our tool. We found the most success when the model was trained on the subset of the TextVQA training data for which the base Qwen2.5-VL-3B-Instruct model gets an average VQA score of $<0.5$ from an 8-shot evaluation on the training set. This evaluation was carried out at a temperature of 1.0 and with at most 10 tokens generated per query.
We originally explored training on the full TextVQA task, but found that tool use was not necessary for a large number of the images. We hypothesize that our base model was trained on this task already, although the exact data mix is not specified in [2]. Additionally, we observe that the model could directly read the text from the full image at the downsized resolution. Based on these results, we attempted to train on only the most difficult images in the training set, where the base model gets an average VQA score of 00 from an 8-shot evaluation. This results in effective learning to use zoom, even in the absence of extrinsic tool use reward, but suffers from poor generalization. These results drove our final mix described above – we overrepresent difficult examples while maintaining some easier ones representative of the full task distribution.
3 Experiments
Table 1: Results on evaluation benchmarks.
Model $\text{TextVQA}_{\text{val}}$ $V^{*}$ Bench HR-Bench 4K HR-Bench 8K
VQA Score Attr Spatial Overall FSP FCP Overall FSP FCP Overall
Closed-source VLMs
Qwen-VL-max [1] 61.5 - - - 65.0 52.0 58.5 54.0 51.0 52.5
GPT4-o [9] 77.4 - - 66.0 70.0 48.0 59.0 62.0 49.0 55.5
Visual Search
SEAL [29] - 74.8 76.3 75.4 - - - - - -
$\text{DC}^{2}$ [27] - - - - 53.0 47.0 50.0 40.5 45.0 42.3
ZoomEye [20] - 93.9 85.5 90.6 84.3 55.0 69.6 88.5 50.0 69.3
Open-source VLMs
LLaVA-HR-X-7B [15] 67.1 51.3 64.5 56.5 57.8 46.3 52.0 42.0 41.3 41.6
Qwen2.5-VL-7B [2] 84.9 80.9 76.3 79.1 85.2 52.2 68.8 78.8 51.8 65.3
Qwen2.5-VL-3B [2] 79.3 81.3 63.2 74.4 81.8 48.5 65.1 80.5 47.3 63.8
Ours 73.4 82.4 76.3 80.1 81.8 48.8 65.3 69.5 46.8 58.1
$\Delta$ v.s. Qwen2.5-VL-3B -5.9 +1.1 +13.1 +5.7 +0.0 +0.3 +0.2 -11.0 -0.5 -5.7
3.1 Experimental Setup
Evaluation benchmarks
We evaluate our models against baselines on both in-domain and out-of-domain high resolution benchmarks. The in-domain reasoning capabilities are evaluated on the validation set of TextVQA. The out-of-domain benchmarks are $V*$Bench [29] and HR-bench [27], used to assess the models’ generalization capabilities.
Baselines
We compare our model to three categories of VLMs: 1) Proprietary VLMs, including Qwen-VL-Max [1] and GPT4-o [9]; 2) visual search methods that engineered for high-resolution tasks, namely SEAL [29], $\text{DC}^{2}$ [27], and ZoomEye [20]; 3) open-source VLMs, such as LLaVA-HR-X-7B [35] and Qwen2.5-VL-Instruct [2] series models.
Training details
We run our training on 4 A100 80GB GPUs. To control the training memory footprint, we limit the rollouts to 3 and the maximum tool calls to 1, while the global batch size is set to 4. We train for 800 steps with a maximum of 2 images per prompt, a completion length of 2048, Adam with $\beta_{2}=0.98$, a cosine learning rate at $61e-$ 6 with 10 warm-up steps, a maximum grad norm of $1.0$, no KL regularization and a clip-higher strategy [33] with $\epsilon_{low}=0.2$ and $\epsilon_{high}=0.28$.
3.2 Results
RL substantially improves VLMs’ performance on some high-resolution tasks
As shown in Table 1, compared to the Qwen2.5-VL-3B baseline, our model achieves 5.7% overall accuracy improvement on $V*$Bench. At the same time, the 3B model trained with GRPO achieves similar performance to the 7B base model, demonstrating greater parameter efficiency given the comparable end-task accuracy. However, our approach decreases in-domain performance on the validation set of TextVQA and does not improve results on HR-Bench. We will discuss these limitations in section 3.3.
End-to-end RL is parameter-efficient compared to visual search methods
Compared to some heavily engineered visual search methods, such as SEAL, our GRPO-trained VLM achieves a similar end-task performance on the $V^{*}$ench. Note that SEAL uses Vicuna-7B as the LLM backbone, whereas our model has only 3 billion parameters, demonstrating superior inference efficiency. Our tool use formulation also results in less complexity at inference time: where visual search methods rely on hierarchical search strategies on images, our method calls the zoom tool to the specific relevant ROI only once.
Low-resolution training does not generalize to 4K & 8K images at inference time
Results in Table 1 show that the performance of the model trained after GRPO is indifferent to the base model, suggesting that such performance gains shown on $V^{*}$ench are not achieved on higher-resolution tasks, namely the HR-bench 4K and 8K tasks. We conjecture that the resolution gap between training and inference makes RL generalize poorly, and our future work will focus on curating training data that is beneficial for models to perform well on ultra-high-resolution tasks.
3.3 Discussion and analysis
We analyze our results in more depth and discuss the contributions of design choices in this section.
Effect of structured reward
Figure 4 shows how the structured reward creates an advantage for successful tool use during training. We plot the log of the average advantage for the following three categories of responses: 1) no tool use; 2) successful tool use; 3) failed tool use. Our training paradigm ensures that successful tool-use calls will be reliably rewarded as training progresses. Moreover, we find that responses without any tool use will receive a negative advantage compared to failed tool-use calls. This supports our belief that our reward structure contains a good signal for the model to tackle tasks that require understanding of visual details in high-resolution images.
Figure 3: Accumulated GRPO advantage for policy trajectories with successful tool use (orange), failed tool use or no tool use during a training run.
Figure 4: Overall accuracy of VLMs on $V^{*}$ench with varied input image resolutions.
Effect of data mix choice
Before using the training examples where the base model achieves $<$ 0.5 VQA score from 8-shot evaluation, we also explored training the model with difficult data only, where only the examples for which the base model gets an average of 0 VQA score are selected, as detailed in 2.1. We find that this data selection strategy hurts rather than helping the model’s performance, resulting in 71.7% accuracy on $V^{*}$ench, which is lower than the baseline. We conjecture that the training model generates too few responses that receive rewards other than for format. Therefore, the training fails to provide any signal for the model to answer those tough questions correctly.
Effect of inference time image resolution
As shown in Figure 4, we notice that using RL to train the VLM for tool-use substantially improves the model’s performance on $V^{*}$ench, even if image resolution is reduced at inference time. This implies that a potential application of our model is to perform low-resolution inference and use a zoom-in tool on high-resolution tasks, thereby improving the system’s inference efficiency. Furthermore, we also notice that resizing the crop dimension to match the input image plays an essential role in accuracy improvements. Without applying this technique, the model’s performance degrades to baseline level.
4 Related Work
4.1 Reasoning with VLMs
As VLMs naturally inherit the reasoning capabilities from their LLM backbone, researchers attempt to teach VLMs with chain-of-thought reasoning traces to ask the model to think before answer [14, 2]. Since the breakthrough of Deepseek-r1 [19, 6], researchers have been focusing on bringing the success of r1-like GRPO training paradigms to VLMs to improve models’ visual reasoning performance. Recent research found that VLMs can achieve superior reasoning capabilities in vision-intensive mathematical tasks through GRPO [8, 34, 21, 30, 4, 26]. However, existing work often requires expensive human annotation or distillation from larger models for generalization. In the meantime, limited effort has been spent on improving VLMs’ vision-oriented reasoning capabilities, where the model can leverage task-specific details residing in images.
4.2 Fine-grained Processing of VLMs
Recent research has pointed out that VLMs often neglect detailed information in images [5], and have blind faith in textual modality [18, 3]. To address this issue, inspired by humans’ behavior in recognizing information from images [32], a series of visual search-based improvements on VLMs have been proposed. Even though such methods can inject additional information of salient regions in images to the VLM, they rely on extrinsic visual search components with complex searching algorithms [29, 20, 11], thus limiting the efficiency of the system. Meanwhile, another line of research attempts to convert the input image into multi-resolution crops and feed them to the VLM end-to-end [23, 13, 35, 27]. Although these methods successfully provide the VLM with image details to tackle downstream tasks, there is computational overhead introduced by the image tokens of the cropped patches. At the same time, these solutions crop the original image generally without recognizing the importance of specific regions in the image to tackle downstream tasks, which may limit their ability to help VLMs solve real-world problems.
4.3 Multimodal Tool use with LLMs
Researchers have been attempting to enable language models to do multimodal tasks by designing language model tool-use systems, i.e., developing systems that treat language models as black boxes to call APIs or use various tools [12, 16]. Without training multimodal foundational models from scratch, previous research has found that iteratively prompting LLMs with vision foundation models can successfully perform numerous vision-language tasks [31, 28, 7]. However, there is only limited research into applying the same approach to vision language models to enhance their vision-oriented reasoning capabilities [25]. A number of researchers have recently applied reinforcement learning to LLMs, including with successful tool-use learning [10, 17]. These efforts have been primarily limited to text-only models, leaving much to be explored in the multimodal domain.
The ideas and results presented in this paper were developed independently and concurrently with those of DeepEyes [36], which corroborate some of our findings. However, our focus is primarily on efficient understanding of visual detail with tool use under constrained resources.
5 Conclusion
We have presented a method for efficiently training small VLMs to understand visual details via GRPO-based reinforcement learning with an option to call external tools, namely to zoom into a region of interest on an image for a closer look. We have shown that this approach allows VLMs to see finer detail where it matters, thereby improving performance on some high-resolution visual understanding benchmarks compared to baselines on the same parameter scale, and even some larger engineered visual search architectures. We have also given a recipe for success involving the right data mix, reward structure, tool interface, and allocation of tokens to external tool results.
Appendix A System and user prompts
We used the following system prompt:
You may call any of the tools exactly one time. You have access to the
following tools to help solve problems:
{tool_descriptions}
For each step:
1. Start by thinking through your reasoning inside <think> tags. Then
either return your answer inside <answer> tags, or use a tool inside
<tool> tags.
2. If needed, use a tool by writing its arguments inside <tool> tags.
Use one line for each argument in the format ’key: value’. The first
line must be ’name: <tool_name>’.
3. You will see the tool’s output inside <result> tags.
4. Continue until you can give the final answer inside <answer> tags.
Tools expect specific arguments. Follow the examples carefully for the
required keys and expected value formats.
Do not make up tools or arguments that aren’t listed.
We used the following user prompt:
The image size is {image_size}. Please thoroughly think through the question and refine your answer while thinking. You should try to collect the visual evidence you need to support your answer. Then, provide your answer. The answer (which you will provide in the <answer> </answer> tags) should be a single word or phrase directly answering the question. Question: {question}